# The Power of Linear Programming in Sponsored Listings Ranking: Evidence from Field Experiments

### v2

@@ Abstract
Sponsored listing is one of the major revenue sources for many prominent online marketplaces, such as Amazon, Walmart, and Alibaba. When consumers visit a marketplace’s webpage for a specific item, in addition to that item, the marketplace might also display a ranked listing of sponsored items from various third-party sellers. These sellers are charged an advertisement fee if a user purchases any of the sponsored items from this listing. Determining how to rank these sponsored items for each incoming visit is a crucial challenge for online marketplaces, a problem known as sponsored listings ranking (SLR). The major difficulty of SLR lies in balancing the trade-off between maximizing the overall revenue and recommending high-quality and relevant ranked listings. While a more relevant ranking may result in more purchases and consumer engagement, the marketplace also needs to take account of the potential revenue when making ranking decisions. Due to the latency requirement and historical reasons, many online marketplaces use score-based ranking algorithms for SLR optimization. Alternatively, recent research also discusses obtaining the ranking by solving linear programming (LP). In this paper, we collaborated with a leading online global marketplace and conducted a series of field experiments to compare the performance of the score-based ranking algorithms and the LP-based algorithms. The field experiment lasted for $19$ days, which included $329.3$ million visits in total. We observed that the LP-based approach improved all major metrics by $1.80\%$ of revenue, $1.55\%$ of purchase, and $1.39\%$ of the gross merchandise value (GMV), compared to an extremely-tuned score-based algorithm that was previously used in production by the marketplace. Our heterogeneous treatment effects analysis revealed that the improved performance stemmed from the LP-based algorithm’s ability to deliver more relevant items to consumers. To enable the practice of the LP-based approaches, we developed an efficient dual-based algorithm for obtaining the ranking within the strict latency requirement. Furthermore, we investigated the flexibility of the LP-based approach to model long-term goals for the marketplace by overall planning. The proposed LP-based model has been launched on Marketplace A’s platform since the beginning of 2023, significantly improving the performance of sponsored listing in multiple dimensions.
@@ 1 Introduction
Sponsored listing is also called “sponsored products” or simply “Ads”, which is a mechanism where third-party sellers allocate additional commission fees to online marketplaces, thereby securing a higher likelihood of their items or advertisements being prominently displayed to consumers. For many online marketplaces, such as Amazon, Walmart, and Alibaba, sponsored listing is one of the major revenue sources. For example, Amazon and Google reported $37.7$ billion and $224.47$ billion of sponsored revenue in 2022 respectively  [^1] https://ir.aboutamazon.com/annual-reports-proxies-and-shareholder-letters/default.aspx, https://abc.xyz/investor/ue.
While the proliferation of sponsored listings can bolster marketplace revenues, the alignment of these prominently displayed items with consumer interests is not guaranteed. In general, displaying more sponsored listings to consumers would simulate revenue for the marketplace, but these exhibited items are not necessarily what consumers are essentially interested in. Thus, for these marketplaces, the critical problem is deciding which items to display and how to rank the sponsored listings to attentively engage in the trade-off between the overall revenue and delivering high-quality, relevant listings for consumers in the meantime. Such a listing ranking problem is called a sponsored listings ranking (SLR) problem, which is heavily studied within the field of recommendation systems.
The design of sponsored listings ranking algorithms in practice involves the following major considerations:
  • Algorithm efficiency: The algorithms must be adequately efficient to meet the real-time display request latency requirement, namely, producing the ranking within at most $0.1$ second.
  • Interpretability and tractability: The ideal algorithm should offer clear and rational objectives and constraints, ensuring that the underlying decision-making process is both interpretable and traceable.
  • Flexibility: The proposed framework should be adaptable, allowing for adjustments in objectives and requirements to accommodate customized SLR challenges in diverse scenarios.
A prevalent industry solution to the SLR problem follows the “predict-then-optimize” paradigm, which comprises the following two steps. First, the “predict” phase involves estimating the conversion rate, which denotes the proportion of consumers completing a desired action, such as a “click” or “purchase.” These actions correspond to the business models of Cost-Per-Click (CPC) and Cost-Per-Acquisition (CPA). After estimating the action rate, the “optimization” phase focuses on selecting the sponsored items for display. This is the focus of our paper.
@@
A common solution widely used in the industry for the “optimization” step is using score-based ranking algorithms, which involves assigning scores to items based on predetermined criteria and then ranking them for display according to these scores. Such algorithms are designed to meet the stringent latency requirements of online platforms, typically generating rankings within 0.1 seconds. The scoring criteria often encompass multiple objectives, including but not limited to the consumer action rate, the relevance of the items to the consumer, and potential revenue generation. A common practice is to employ a score function that linearly combines multiple objectives that fit online marketplaces’ interests the most (Liu et al.,, 2015; Cheng et al.,, 2016; Covington et al.,, 2016; Okura et al.,, 2017; Zhou et al.,, 2018).
However, there are several inherent limitations associated with the score-based algorithm. First, the quality of the final ranking entirely depends on the definition of the score function, posing a significant challenge for online marketplaces in picking the most suitable score function. The obtained ranking might vary greatly with the choice of score functions, and for those that are not direct measurements of the key economic features, the ranking results might be sub-optimal and unable to achieve Pareto efficiency. Furthermore, through the multi-objective score function, it is difficult to explicitly guarantee individual objectives’ performance. More specifically, the score-based method does not incorporate scenarios where marketplaces would like to set constraints on resources and considerations such as inventory, sales performances, and fairness. To resolve these issues, we propose and study an alternative approach called the LP-based algorithm in this paper. Specifically, a ranking problem can be formulated as a constraint mixed integer programming (MIP) problem. Note that since MIP is NP-hard, we relax it to an LP problem for tractability. Unfortunately, even with state-of-the-art LP solvers, solving the LP directly still cannot meet the strict latency requirement for online marketplaces. With these insights, in this paper, we first propose a dual-based algorithm for approximately solving the MIP formulation of the SLR problem that can find an approximate solution within $0.1$ second to meet the latency requirement. Then, we present a detailed field study, in collaboration with a leading online marketplace’s platform (referred to as “Marketplace A” for proprietary consideration), that empirically compares the performance of the proposed LP-based algorithm and an industry-level extremely-tuned score-based algorithm. The empirical study clearly showcases the effectiveness of the LP-based algorithm, which improves three key metrics for Marketplace A, namely, purchase, gross merchandise value, and marketplace revenue, in a 19-day A/B test that was performed at the online marketplace in April 2022. Subsequently, the LP-based algorithm was put into production at Marketplace A in January 2023 to help the marketplace provide high-valued sponsored items to consumers.
The contributions of the paper can be summarised as follows:
The major contribution of the paper is an extensive field study (A/B test) on Marketplace A comparing the performance of the LP-based algorithm and the extremely tuned score-based algorithm previously used in production at Marketplace A. Conducted over a span of 19 days in April 2022, the field study encompassed approximately 329 million impressions and yielded conclusive evidence of the LP-based algorithm’s superior performance, as manifested in a 1.80% increment in revenue, a 1.55% augmentation in purchase rates, and a 1.39% elevation in Gross Merchandise Value (GMV). We further present a careful performance analysis of the experiment results, which showcases that (i) the LP-based algorithm displays more relevant items to consumers, i.e., items that fit consumers’ interests, thus resulting in more sales, a higher conversion rate, and eventually more revenue; (ii) the average transaction price of the LP-based algorithm is no higher than the score-based algorithm, which demonstrates that the additional revenue does not originate from a lifted item price. To the best of our knowledge, this is the first empirical study comparing the LP-based and score-based approaches for sponsored listings ranking problems, a core problem faced by online marketplaces. After another round of an 18-day A/B test in September 2022, which demonstrated consistent results, the LP-based algorithm was put into production at Marketplace A in January 2023, generating significant growth in GMV, purchase, and revenue for Marketplace A.
Another contribution is the development of an LP-based algorithm for the SLR problem. Starting with the SLR problem formulated as a MIP problem, we apply the LP relaxation to obtain a primal-dual pair problem, then use binary search to recover the optimal dual variable. Through duality, our algorithm can obtain an integral solution that is feasible for the MIP problem, and usually with a minor optimal gap, as shown in section 3.3 through simulations. Additionally, we show that our algorithm is more than $20$ times faster than the state-of-the-art commercial LP solver on SLR instances of typical size with 50 slots and 500 items, enabling LP-based algorithms in practice for the latency requirement.
Last, we present extensions with further modeling power of the LP-based methods that can be used to solve problems with additional constraints in an online setting. In practice, beyond relevance and revenue, online marketplaces often face other constraints: consumers would prefer to browse fewer sponsored items but more items from a variety of sellers; sponsored sellers are concerned with inventory management and sales performances, and wish for their listed items to be displayed to as many consumers as possible. As these constraints exist globally and might vary over time, the score-based algorithms do not have the flexibility to address them. To overcome this limitation, we extend our initial LP-based model by constructing a comprehensive large-scale overall problem that incorporates all local optimization problems for each impression accumulated over a period. Subsequently, we introduce global constraints to the overall problem, which can be solved by the newly-developed large-scale LP solver PDLP (Applegate et al.,, 2021). After solving the overall problem offline and obtaining the optimal dual variables, we then utilize these dual variables in online LP-based SLR problems.
Note that score-based algorithms are the golden standard in the industry for sponsored listing and other ranking problems; this work can provide empirical evidence for the marketplace to implement the LP-based algorithms.
@@ 1.1 Literature reivew
Our work provides an approach for the ranking step in the recommendation system. In a recommendation system, the first step is to estimate the conversion rate of displaying the impression by machine learning algorithms, which has been quite mature and provide reasonable estimations in the online advertising industry (Liu et al.,, 2015; Cheng et al.,, 2016; Covington et al.,, 2016; Okura et al.,, 2017; Zhou et al.,, 2018). After estimating the conversion rates, online platforms are tasked with developing a ranking model that displays candidate items to consumers in a designated order, which is the focus of this paper. A majority of previous studies focus on the score-based ranking model (Zehlike et al.,, 2022; Renda and Straccia,, 2003; Montague and Aslam,, 2001; Manmatha et al.,, 2001), that is, calculating a score for each item through a customized scoring function, and then selecting the displayed items by ranking the scores. While the ranking algorithms are similar among various studies, the scoring function, however, might vary, depending on the objectives of each online platform. For instance, Covington et al., (2016) deployed a deep neural network model and used logistic regressions to generate scores for each video content on YouTube. Google Play (Cheng et al.,, 2016), instead, used only the estimated conversion rate as the score for each item, therefore, aiming to recommend the most relevant items to consumers. Netflix (Amatriain,, 2013) adopted a score function that linearly combines popularity and predicted rating for each video. As we will introduce more details in 2, Marketplace A also utilizes a score-based ranking method whose objective combines different sources of revenue (Ge et al.,, 2020). In the case of “cold start”, i.e., the launch of a new advertisement/item, where the data for conversation rate estimation is limited, Ye et al., (2023) proposed a bandit algorithm that significantly increased the success rate of cold start in field experiments.
@@
An inevitable drawback of the score-based method is that even though online platforms aim to achieve Pareto efficiency over multi-objective (Lin et al.,, 2019), it is hard to guarantee the performance of individual objectives. An alternative approach is to formulate the problem as a mathematical programming problem, that is, maximizing an objective function while setting constraints on some other metrics and resources. Such a problem can be formulated as a mixed integer programming (MIP) problem (Celis et al.,, 2017), and sometimes referred to as an advertisement allocation problem (Mehta et al.,, 2013). We can also associate this problem with some well-studied literature on assignment problem (Ross and Soland,, 1975) and knapsack problem (Salkin and De Kluyver,, 1975). However, as the number of constraints increases, the MIP also becomes increasingly harder to solve, requiring various approximation techniques. A more comprehensive review can be referred to (Cattrysse and Van Wassenhove,, 1992; Pentico,, 2007; Karp et al.,, 1990). Particularly, Celis et al., (2017) considered the ranking problem with fairness constraints. Based on LP relaxation, they recovered integral solutions with small constraint violations. Asadpour et al., (2023) studied a ranking model that aims to maximize consumer engagement for online retail, proposing an algorithm by LP relaxation and randomized rounding. Further, through LP-relaxation and primal-dual pairs, Google (Bhalgat et al.,, 2012) addressed an advertisement allocation problem with nested packing constraints, and then approximated it with an LP-based primal-dual algorithm; Microsoft (Chen et al.,, 2011) proposed a bidding algorithm for advertisement allocation with budget constraints. Inspired by (Chen et al.,, 2011), Alibaba (Zhang et al.,, 2018)) developed a similar algorithm for a global advertisement allocation problem that optimized over accumulated search queries. It is noted that, despite the availability of standard algorithms in the existing literature for advertisement allocation problems, the introduction of new problems with distinct objectives or constraints necessitates the customization and adjustment of these paradigms. Our algorithm belongs to this category, and we provide an efficient approximate algorithm for solving the MIP problem and the first field study comparing the performance of scored-based algorithm and LP-based algorithms.
@@
2 Sponsored listing and score-based ranking at Marketplace A
In this section, we present an overview of the marketplace of Marketplace A, with a focus on sponsored listing. Then, we discuss the score-based ranking algorithm previously deployed at Marketplace A.
2.1 An overview of Marketplace A’s Marketplace And sponsored listing
The essence of the marketplace is to match the search requests from consumers by providing them with a carefully selected range of items listed by sellers, thereby enabling transactions. Providing a carefully ranked list of items is referred to as the ranking problem, which is a major task for the online marketplaces. Generally speaking, Marketplace A’s marketplace is tasked with two key ranking problems, namely, the keyword ranking (KR) and the sponsored listings Ranking (SLR). The KR pertains to the problem through which the search engine displays items according to consumer-entered keywords, which is a problem shared by not only online marketplaces but also many primary search engines like Google and Bing. For instance, when a consumer visits Marketplace A’s webpage and searches for a keyword, a list of ranked items will be displayed correspondingly as exhibited in Figure 1(a). The study of how to rank these items is called KR. Then, the consumer may click one of the displayed items that she is interested in, and enter its main page (Figure 1(b)). This specific item is called seed item of the page. Below the seed item is a carefully selected list of related items, which is called Sponsored listings items. The objective of SLR is to recommend consumer-related alternatives or complementary items that may pique the interest of consumers who might not be captivated by seed item. Sponsored listings is also phrased as “sponsored items”, “often bought together”, “you might also like”, across different online marketplaces such as Amazon and BestBuy.
Marketplace A has multiple revenue sources, among which the sponsored listing is one of the most significant. In this paper, we specialize in SLR. Particularly, the goal of SLR is to display relevant items for consumers and generate sizable revenue. For sponsored listings, Marketplace A’s revenue primarily comes from two components: 1) take fee (also called organic revenue) and 2) sponsorship fee (also called Ad revenue). For every item sold at the marketplace, a certain percentage of the price is taken by Marketplace A as a commission fee, which is referred to as take fee. The percentage (called take rate) is tiered and ranges from 3% to 15%, depending on the item’s category and price. In addition, the sellers can voluntarily sponsor their items by contributing additional fees to boost the visibility of items to consumers. Such fees will be charged proportionally to items’ prices, where the proportion is referred to as Ad Rate.
From the historical data of Marketplace A, we observed that organic revenue is highly correlated with purchase count, as a higher purchase-through rate (PTR) will lead to more purchases and thus a higher organic revenue. It means that more relevant rankings would lead to more organic revenue and less Ad revenue. Contrariwise, a new policy that generates revenue improvement usually comes with more Ad revenue and the curse of less organic revenue. Spontaneously, the trade-off between relevance and revenue could also be viewed as the one between organic revenue and Ad revenue.
When a consumer opens a webpage with a seed item, Marketplace A first selects a collection of qualified candidate items. From now on, we use the set $\mathcal{N}=\{1,2,...,n\}$ to denote this collection. On the other hand, depending on the display settings across various platforms (Figure 2), the slot number $m$ might differ as well (up to $50$.) Then, Marketplace A is tasked to (1) choose a subset of $\mathcal{N}$ according to the slot number $m(m\leq n)$ and then (2) display this subset of items in ranked order. The consumer may navigate to a promoted item and make the purchase. While looking after the welfare of all three parties (consumers, sellers, and the marketplace), the objective of SLR is to display the ranking that not only fits consumers’ interests but is also able to generate substantial revenue for both the sellers and the marketplace.
(a) Search Listings display
(b) Sponsored listings display
Figure 1: Display settings of Marketplace A’s marketplace for keyword Ranking (left) and sponsored listings ranking (right)
(a) Sponsored listings display for iPhone
(b) Sponsored listings display for iPad
Figure 2: Sponsored listings display settings for iPhone (left) and iPad (right)
2.2 Score-based ranking for sponsored listing at Marketplace A

### v3

# The Power of Linear Programming in Sponsored Listings Ranking: Evidence from a Large-Scale Field Experiment

@@ Abstract
Sponsored product advertisements constitute a major revenue source for online marketplaces such as Amazon, Walmart, and Alibaba. A key operational challenge in these systems lies in the Sponsored Listings Ranking (SLR) problem, that is, determining which items to include and how to rank them to balance short-term revenue with long-term relevance and user experience. Industry practice predominantly relies on score-based algorithms, which construct heuristic composite scores to rank items efficiently within strict real-time latency constraints. However, such methods offer limited control over objective trade-offs and cannot readily accommodate additional operational constraints. We propose and evaluate a Linear Programming (LP)-based algorithm as a principled alternative to score-based approaches. We first formulate the SLR problem as a constrained mixed integer programming (MIP) model and develop a dual-based algorithm that approximately solves its LP relaxation within 0.1 second, satisfying production-level latency requirements. In collaboration with a leading online marketplace, we conduct a 19-day field experiment encompassing approximately 329 million impressions. The LP-based algorithm significantly outperforms the industry-standard benchmark in key marketplace metrics, demonstrating both higher revenue and maintained relevance. Mechanism analyses reveal that the performance gains are most pronounced when the revenue–relevance tradeoff is stronger. Our framework also generalizes to settings with inventory, sales, or fairness constraints, offering a flexible and deployable optimization paradigm. The LP-based algorithm was deployed in production at our partner marketplace in January 2023, marking a rare large-scale implementation of a mathematically grounded ranking algorithm in real-world online advertising.
Keywords: online advertising, sponsored listings rankings, linear programming, latency, revenue-relevance tradeoff, field experiment
@@ 1 Introduction
Sponsored product advertisements constitute a primary revenue stream for most major online marketplaces, including Amazon, Walmart, and Alibaba. For instance, Amazon and Google reported \$37.7 billion and \$224.47 billion in sponsored advertising revenues in 2022, respectively. [^1] https://ir.aboutamazon.com/annual-reports-proxies-and-shareholder-letters/default.aspx ; https://abc.xyz/investor/ue In such advertising systems, third-party sellers pay additional commission fees to online marketplaces to increase the visibility of their products, thereby enhancing the likelihood that their items are prominently displayed to consumers.
A particularly prevalent format of sponsored product advertising is the sponsored product listing, which is typically displayed under a seed product. In this format, promoted items are interleaved with organically ranked products and may appear in sections such as “Sponsored Related Items,” “Products Related to This Item,” or “You May Also Like,” with platform-specific variations in terminology. Figure 1 provides an illustrative example from Amazon. Within these listings, both sponsored and non-sponsored products coexist without predetermined slots reserved for sponsorships. This intermixing introduces substantial complexity to determining which items to include in the listings and how to optimally rank them, a problem we refer to as the Sponsored Listings Ranking (SLR) problem, which is the primary focus of this study.
Figure 1: Sponsored Product Listings Example
The primary challenge in SLR design lies in managing the trade-off between revenue generation and relevance. While displaying more sponsored products can increase short-term platform revenue, such placements may not align with consumer preferences and could degrade user experience. Hence, the platform must balance two competing objectives: maximizing revenue while maintaining the relevance and quality of displayed items to sustain long-term consumer engagement. Achieving this balance requires a principled approach to ranking.
Beyond this fundamental trade-off, the practical design of SLR algorithms involves several additional considerations. First, in terms of computational efficiency, the algorithm must satisfy stringent latency requirements, typically producing a ranked list within 0.1 seconds to meet real-time display constraints. Second, with respect to interpretability and tractability, the ranking framework should embody clear objectives and explicit constraints, allowing the decision-making process to remain transparent and auditable. Third, the algorithm must be flexible and adaptable, enabling adjustments in objectives, constraints, or weighting schemes to accommodate platform-specific business goals and varying marketplace contexts.
To address practical considerations, particularly the stringent latency requirement, most industry solutions to the SLR problem adopt a “predict-then-optimize” paradigm comprising two stages. In the prediction stage, platforms estimate user action rates, such as click-through rate (CTR) or purchase-through rate (PTR), which correspond to the prevailing Cost-Per-Click (CPC) and Cost-Per-Acquisition (CPA) business models. In the optimization stage, the system determines which sponsored items to display and how to rank them. Our proposed algorithmic solution intervenes at this optimization stage while keeping the prediction stage identical to that used in industry.
@@
A widely adopted approach for the optimization stage in practice is the score-based ranking algorithm, which assigns each item a composite score based on predefined criteria and ranks items according to these scores. These algorithms are favored for their computational efficiency, typically producing rankings within 0.1 seconds to satisfy real-time latency requirements. The scoring criteria generally integrate multiple objectives, such as predicted consumer action rate, item relevance, and potential revenue generation. In most implementations, platforms employ a linear combination of these objectives to construct the score function that best aligns with their business priorities (Liu et al.,, 2015; Cheng et al.,, 2016; Covington et al.,, 2016; Okura et al.,, 2017; Zhou et al.,, 2018).
However, score-based algorithms face several inherent limitations. First, the final ranking quality depends critically on how the score function is specified, making it difficult for marketplaces to select an appropriate formulation. Rankings may vary substantially across different scoring functions, and when the chosen function does not directly represent the underlying economic objectives, the results can be suboptimal and fail to achieve Pareto efficiency. Moreover, aggregating multiple objectives into a single composite score provides no explicit control over the performance of individual objectives. In particular, this approach cannot readily accommodate operational constraints that marketplaces may wish to impose, such as inventory limits, sales targets, or fairness considerations.
To address these challenges, we propose and examine an alternative approach, referred to as the Linear Programming (LP)-based algorithm, in this paper. Specifically, we formulate the ranking problem as a constrained mixed integer programming (MIP) model. Because solving the MIP exactly is NP-hard, we relax it to a tractable LP formulation. However, even with state-of-the-art solvers, directly solving the LP remains too computationally intensive to satisfy the stringent latency requirements of online marketplaces. Motivated by these constraints, we develop a dual-based algorithm that approximately solves the original MIP formulation of the SLR problem within 0.1 second, thereby meeting the platform’s real-time demands. We then validate the proposed approach through a large-scale field experiment conducted in collaboration with a leading online marketplace (referred to as Marketplace A for confidentiality). The 19-day experiment, run in April 2022 and encompassing approximately 329 million impressions, demonstrates that the LP-based algorithm substantially outperforms the highly tuned industry-standard score-based algorithm, leading to significant improvements in key marketplace performance metrics.
We further examine the underlying mechanism through which our LP-based algorithm outperforms the score-based benchmark. The key difference lies in how the revenue–relevance tradeoff is modeled. In the score-based algorithm, the industry convention is to frame this tradeoff as one between organic revenue and ad revenue (see Section 2 for details on the revenue structure). In contrast, our LP-based algorithm adopts a more disciplined formulation of the revenue–relevance tradeoff: rather than separating revenue by source, it maximizes total marketplace revenue subject to a minimum relevance constraint. Empirically, we provide evidence supporting this mechanism. The performance gains of the LP-based algorithm are more pronounced precisely when the revenue–relevance tradeoff is stronger. Furthermore, the algorithm demonstrates advantages in both the item selection and conditional ranking stages, consistently identifying items with higher expected revenue.
Our key contribution lies in developing a new ranking algorithm, the LP-based algorithm, for the SLR problem, which can simultaneously achieve higher revenue and satisfy the stringent latency requirements of production environments. We further provide empirical evidence from a large-scale field experiment demonstrating the superior performance of the proposed approach. To the best of our knowledge, this paper represents the first empirical study to both document large-scale experiment evidence and uncover the mechanisms underlying the LP-based algorithm’s advantage over score-based algorithms. In addition, our LP-based framework has strong modeling flexibility, enabling it to incorporate additional operational constraints, such as inventory limits, sales targets, or fairness considerations, within an online optimization setting.
Our proposed algorithm was deployed in production at Marketplace A in January 2023, where it helps the platform deliver high-value sponsored items to consumers in real time. Notably, score-based algorithms have long served as the industry standard for sponsored listing and other ranking problems. Our work, by contrast, introduces a principled and practically viable alternative that has been successfully implemented in production, demonstrating both theoretical soundness and real-world effectiveness.
@@ 1.1 Literature reivew
Our work provides an approach for the ranking step in the recommendation system. In a recommendation system, the first step is to estimate the conversion rate of displaying the impression by machine learning algorithms, which has been quite mature and provide reasonable estimations in the online advertising industry (Liu et al.,, 2015; Cheng et al.,, 2016; Covington et al.,, 2016; Okura et al.,, 2017; Zhou et al.,, 2018). After estimating the conversion rates, online platforms are tasked with developing a ranking model that displays candidate items to consumers in a designated order, which is the focus of this paper. A majority of previous studies focus on the score-based ranking model (Zehlike et al.,, 2022; Renda and Straccia,, 2003; Montague and Aslam,, 2001; Manmatha et al.,, 2001), that is, calculating a score for each item through a customized scoring function, and then selecting the displayed items by ranking the scores. While the ranking algorithms are similar among various studies, the scoring function, however, might vary, depending on the objectives of each online platform. For instance, Covington et al., (2016) deployed a deep neural network model and used logistic regressions to generate scores for each video content on YouTube. Google Play (Cheng et al.,, 2016), instead, used only the estimated conversion rate as the score for each item, therefore, aiming to recommend the most relevant items to consumers. Netflix (Amatriain,, 2013) adopted a score function that linearly combines popularity and predicted rating for each video. Yang et al., (2019) proposes a linear programming based algorithm for online advertisement bid, without considering the ranking. As we will introduce more details in section˜2, Marketplace A also utilizes a score-based ranking method whose objective combines different sources of revenue (Ge et al.,, 2021). In the case of “cold start”, i.e., the launch of a new advertisement/item, where the data for conversation rate estimation is limited, Ye et al., (2023) proposed a bandit algorithm that significantly increased the success rate of cold start in field experiments.
@@
An inevitable drawback of the score-based method is that even though online platforms aim to achieve Pareto efficiency over multi-objective (Lin et al.,, 2019), it is hard to guarantee the performance of individual objectives. An alternative approach is to formulate the problem as a mathematical programming problem, that is, maximizing an objective function while setting constraints on some other metrics and resources. Such a problem can be formulated as a mixed integer programming (MIP) problem (Celis et al.,, 2017), and sometimes referred to as an advertisement allocation problem (Mehta et al.,, 2013). We can also associate this problem with some well-studied literature on assignment problem (Ross and Soland,, 1975) and knapsack problem (Salkin and De Kluyver,, 1975). However, as the number of constraints increases, the MIP also becomes increasingly harder to solve, requiring various approximation techniques. A more comprehensive review can be referred to (Cattrysse and Van Wassenhove,, 1992; Pentico,, 2007; Karp et al.,, 1990). Particularly, Celis et al., (2017) considered the ranking problem with fairness constraints. Based on LP relaxation, they recovered integral solutions with small constraint violations. Asadpour et al., (2023) studied a ranking model that aims to maximize consumer engagement for online retail, proposing an algorithm by LP relaxation and randomized rounding. Further, through LP-relaxation and primal-dual pairs, Google (Bhalgat et al.,, 2012) addressed an advertisement allocation problem with nested packing constraints, and then approximated it with an LP-based primal-dual algorithm; Microsoft (Chen et al.,, 2011) proposed a bidding algorithm for advertisement allocation with budget constraints. Inspired by (Chen et al.,, 2011), Alibaba (Zhang et al.,, 2018)) developed a similar algorithm for a global advertisement allocation problem that optimized over accumulated search queries.
It is worth noting that although standard algorithms for advertisement allocation problems are well established in the existing literature, newly formulated problems with distinct objectives or constraints often require tailored adaptations of these paradigms. Our proposed algorithm falls into this category. Specifically, we develop an efficient approximate algorithm to solve the underlying mixed integer programming (MIP) formulation, with direct application to the SLR problem. Furthermore, we present the first field study that empirically compares the performance of score-based and LP-based algorithms.
In the context of recommender systems, the primary design objective is to ensure low latency, that is, the ability to generate rankings within a strictly limited response time. Consequently, unlike assortment optimization problems (which we will discuss shortly), ranking problems typically do not treat item interactions as a central concern. Most studies therefore assume that position weights (often referred to as “position bias”) are independent of the item assortment (Chen et al.,, 2023; Agarwal et al.,, 2019; Hofmann et al.,, 2013) and develop methods such as propensity-score-based estimation to correct for this bias (Chen et al.,, 2023; Hofmann et al.,, 2013; Joachims et al.,, 2017).
@@
2 Institutional Context and Score-Based Ranking
We begin this section by outlining the relevant institutional details of our partner marketplace, including how sponsored listings rankings and the advertising system operate in Marketplace A. We then describe the current score-based ranking algorithm employed by Marketplace A.
2.1 Sponsored Listings Rankings (SLR)
Sponsored product advertisements are a major source of revenue for most online marketplaces, including Amazon, Alibaba, and our partner marketplace. These sponsored products typically appear in two key locations where they are relevant as a ranking problem, which we define as generating a carefully ordered list that includes both sponsored and non-sponsored products.
The first location is the keyword search results page, where the ranking process is referred as Keyword Ranking (KR) in our partner marketplace. KR addresses the challenge of displaying and ranking items based on consumer-entered search queries. For example, when a consumer searches for a keyword on Marketplace A, a ranked list of relevant items is presented, as illustrated in Figure 2(a). The task of determining the order of these listings constitutes the KR problem.
The second common location for sponsored products is the product detail page, which consumers reach after clicking on a specific product. In this context, sponsored products may appear in sections such as "Sponsored Related Items," "Products Related to This Item," or "You May Also Like," with specific terminology varying across platforms. The corresponding display format in Marketplace A is shown in Figure 2(b). The ranking of items in this section, which includes both sponsored and non-sponsored products, is referred to as Sponsored Listings Ranking (SLR) in our partner marketplace and is the primary focus of our study. The main product featured on the page is typically referred to as the seed item. As with other marketplaces, Marketplace A notifies users that the listings may include sponsored items, but does not disclose which specific items are sponsored.
(a) Keyword Ranking
(b) Sponsored Listings Ranking
Figure 2: Example of Two Ranking Problems in Marketplace A
In Marketplace A, as in many similar online marketplaces, the goal of SLR is twofold: to display relevant items to consumers and to generate substantial revenue. Achieving this dual objective is a central challenge not only for Marketplace A but also for many of its industry peers. On the one hand, Marketplace A aims to maximize revenue, with the revenue structure detailed in Section 2.2. On the other hand, the platform places high importance on maintaining a positive consumer shopping experience, which is crucial for its long-term growth. To support this, more relevant items, defined as those closely related to the seed item, should be ranked higher in the sponsored listings. A detailed explanation of how relevance is measured is provided in Section 2.3. Balancing this revenue-relevance tradeoff is the core task of the ranking phase, particularly under the predict-then-optimize framework adopted by Marketplace A and widely used across the industry.
Beyond practical constraints, especially the stringent latency requirement, three inherent characteristics of the SLR problem make ranking a particularly critical and challenging task, similar to those encountered in recommender systems. First, the number of sponsored items displayed is not fixed. As discussed in Section 2.3, it can range from zero to the total number of available slots, entirely determined by the ranking algorithm. Second, there are no fixed display positions reserved for sponsored items; their placement is dynamically assigned based on the ranking algorithm’s output. Third, from the consumer’s perspective, sponsored items are not marked with any prominent label, making them indistinguishable from organic listings. These features contrast sharply with traditional sponsored ads in keyword search engines like Google. In those systems, ranking is relatively less central: a small number of fixed, clearly labeled ad slots are shown, and platforms can rely on simple boosting mechanisms for selected sponsored items without needing to rank all results across pages. As we will discuss in Section 6, our KR setup also differs from that of Google, which allows our proposed ranking algorithm to be extended to the KR setting in our context.
In Marketplace A, KR and SLR are managed by separate teams within distinct organizational units that operate independently, a common practice in large-scale marketplaces. As a result, a product that appeared in the KR results but was not clicked by consumers may still appear in the SLR section. There is currently no mechanism in place to prevent this duplication, highlighting a broader challenge faced by large organizations that rely on algorithmic optimization across siloed teams. All empirical results reported in Section 5 are based on the SLR. In addition, we provide empirical evidence on how changes to the SLR can influence outcomes in the KR in Section 5.3.
2.2 Revenue Structure and Advertising System
Revenue from sponsored listings comprises two components: (1) the take fee (also referred to as organic revenue) and (2) the sponsorship fee (also referred to as ad revenue). For every item sold on the platform, Marketplace A collects a commission, called the take fee, which is a fixed percentage of the item’s final sale price. This percentage, known as the take rate, is tiered and ranges from 3% to 15%, depending on the item’s category and price.
In addition, sellers may choose to voluntarily sponsor their items by paying an additional fee to boost visibility. This fee is calculated as a percentage of the final sale price and is referred to as the ad rate. Marketplace A employs a bidding system for setting the ad rate: sellers specify a bid as a percentage of the sale price when setting up a sponsored product, with allowable bids ranging from 1% to 100%. While sellers have the flexibility to choose any bid within the allowable range, the effectiveness of their sponsored product (i.e., its visibility and ranking) depends on how competitive their bid is relative to the trending rate.
Importantly, Marketplace A adopts a cost-per-acquisition (CPA) model, in which the sponsorship fee is charged only when a product is purchased through a sponsored listing. While our discussion focuses on the SLR setting, it is worth noting that the same revenue structure, including the take rate and ad rate, also applies to the KR setting.
We make two clarifying notes. First, the revenue structure, comprising both organic revenue and advertising revenue, is widely adopted by major marketplaces such as Amazon and Alibaba. Accordingly, the fundamental revenue–relevance tension inherent in the SLR problem is broadly applicable. Second, under this revenue structure, consumers are not constrained to purchase only one item from the sponsored listing. The final goal is to maximize revenue while maintaining relevance and quality of displayed items to sustain long-term consumer engagement.
2.3 Score-Based Ranking